# Phonon assisted light absorption and emission in cubic-Boron Nitride

### v1

@@ Abstract
Cubic boron nitride (cBN) is a wide–bandgap polymorph of boron nitride whose optical response remains only partially understood due to the coexistence of indirect electronic transitions and strong exciton–phonon coupling. Using first-principles many-body perturbation theory, we investigate the optical properties of cBN by combining GW quasiparticle corrections with Bethe-Salpeter equation calculations of excitonic effects. Phonon-assisted absorption and emission processes are explicitly included through the exciton–phonon coupling formalism. We find that phonon–mediated optical transitions provide a dominant contribution to both absorption and luminescence spectra, partially reconciling the discrepancy between the theoretical optical gap ( eV) and experimental emission around 6–7 eV. Our results demonstrate the importance of including exciton–phonon interactions for the correct interpretation of experimental spectra, offering new insights into light emission in wide-bandgap materials.
I Introduction
Boron nitride (BN) exhibits a rich polymorphism, encompassing phases ranging from the common hexagonal (hBN) and cubic (cBN) structures to more complex polytypes.Cazorla and Gould (2019) A significant challenge in the study of these materials is the frequent coexistence of different phases and the difficulty in obtaining perfectly pure samples. This complexity underscores the need for precise characterization techniques. While Raman spectroscopy is a widely used characterization tool in materials science, its utility for BN polymorphs is often limited due to the material’s few active Raman modes and the deep UV band gap forcing nonresonant conditions. Consequently, luminescence spectroscopy emerges as a far more powerful and informative probe. The sensitivity of luminescence to electronic structure, defects, and stacking order makes it exceptionally well-suited for characterizing BN nanostructures, where it can reveal details that remain inaccessible to Raman scatteringJaffrennou et al. (2007). For example, luminescence has proven to be a versatile tool, successfully distinguishing between different stacking sequences in the hexagonal phaseZanfrognini et al. (2023) and probing the effects of strain.Lechifflart et al. (2022); Schué (2017)
@@ I Introduction
Here, we present a first-principles investigation of the optical absorption and luminescence of cBN that explicitly includes exciton–phonon coupling within a many-body perturbation theory framework. By combining GW quasiparticle corrections, Bethe–Salpeter equation calculations, and phonon-assisted optical processes, we assess the role of indirect excitonic transitions in shaping the optical spectra of cBN and clarify their relevance for the interpretation of experimental observations.
@@
This information is critical, given that cBN is a large-gap material where strongly bound excitons are expected and the coupling between excitons and phonons is known to be profound in other BN polymorphs.Schué et al. (2019b); Elias et al. (2021b) Furthermore, understanding exciton–phonon coupling in cBN has broader implications, particularly in light of recent studies on small polaron formation in doped cBN systemsShirodkar et al. (2023) and the reported observation of the electron–phonon Fano effect in cBN.Zhu et al. (2023).
The manuscript is organized as follows: in Sec. II we summarize all theoretical methods used in our calculations including the excitonic properties and the coupling with phonons; Sec. II.3 contains all computational parameters and code references used in this work; in Sec. III we present our results and compare them with the existing literature and experimental measurements; finally, in Sec. IV we draw conclusions and discuss future perspectives for this work.
II Theoretical methods
The ground-state properties of cubic boron nitride are calculated within density functional theory (DFT), as implemented in the Quantum ESPRESSO package Giannozzi et al. (2017). The experimental lattice parameter of 3.615 Å reported in Ref. Knittle et al., 1989 is used throughout. Vibrational properties and electron–phonon matrix elements are obtained using density functional perturbation theory (DFPT) Giannozzi et al. (2017). All parameters that enter the calculation are reported in Sec. II.3. Following our investigation of the ground-state properties, we shifted our focus to exploring the electronic structure and optical properties of BN. To study the electronic bands of BN, we diagonalized the Kohn-Sham (KS) Hamiltonian. However, as DFT is known to have limitations in accurately describing the electronic structure, including underestimating the electronic gapMartin et al. (2016), we used Many-Body Perturbation Theory implemented in the Yambo code Sangalli et al. (2019) to correct the KS eigenvalues. By doing so, we calculated the quasiparticle bandstructure as:
where are the corrected quasiparticle energies, is the renormalization factor, the exchange correlation KS potential and the self-energy is calculated in the GW approximation Martin et al. (2016), at frequency . The screened interaction is calculated using the Godby-Needs plasmon–pole model.Stankovski et al. (2011)
Starting from the quasiparticle bandstructure we proceeded to calculate the optical response function, including the electron–hole interaction. To achieve this, we solved the Bethe-Salpeter Equation (BSE) Strinati (1988), which includes both local–field effects and screened electron–hole attraction. The BSE can be reformulated as an eigenvalue problem, which is defined by a two-particle Hamiltonian Martin et al. (2016); Gatti and Sottile (2013) written on a basis of electron–hole transitions . In this work, we solved the BSE in the Tamm-Dancoff approximation, including therefore only resonant transitions, , where the momentum is in the first Brillouin zone ( and denoting valence and conduction bands, respectively). The matrix elements of the BSE are written in the following manner:
where is the energy associated with the above mentioned transition from valence to conduction, is the Coulomb potential derived from the variation of the Hartree term, and is the screened electron–hole interaction derived from the screened exchange.Strinati (1988) The energies that enter the BSE include the quasiparticle corrections calculated within the approximation. After diagonalization of Eq. 2 we obtain the eigenvectors and eigenvalues that are the eigenstates and the energies of the excitons (labeled by the index ), the stationary state of the BSE. The exciton energies and wave-functions at are used to build up the macroscopic dielectric function as:
@@ II Theoretical methods
where are the excitonic dipoles, are the dipole matrix elements between the Kohn-Sham states and is a small broadening term added to mimic the resolution of experimental spectra. To speed up convergence of the BSE with the number of k-points, we used the double-grid approach described in Ref. Alliati et al., 2022, see Sec II.3 for details.
II.1 Phonon-assisted light absorption
@@
For the phonon-assisted absorption/emission, we follow the approach derived in Ref. Lechifflart et al., 2023 with two modifications. First, the phase of the exciton–phonon matrix elements has been fixed using the code and the approach described in Ref. Nalabothula, 2025. Second, the definition of phonon-assisted transition dipoles has been corrected to be gauge consistent. This corresponds to including also the off-diagonal matrix elements of the exciton-phonon self-energy of Ref. Lechifflart et al., 2023. Then, the first order correction to the dielectric constant Eq. 3 induced by the coupling with phonons can be written as:
@@
here, are the excitonic dipoles and are the phonon-assisted coupling dipole moments. Eq. 4 describes the satellites generated by the emission or absorption of phonons by an exciton, and they appear as a peak at the energy of the finite-momentum excitons plus or minus one phonon energy. The ‘’ signs refer to phonon emission or absorption processes. The appearance of satellites also induces a renormalization of the purely exciton dielectric constant, a term that we do not consider here because the effect is small, and we investigate the two effects separately. For more details, see Ref. Lechifflart et al., 2023; Marini et al., 2024. The phonon-assisted dipole moments can be written explicitly as
where are the exciton–phonon coupling matrix elements at momentum Lechifflart et al. (2023); Marini et al. (2024)(since luminescence is always mediated by the recombination of an intermediate optical exciton, depends on a single momentum corresponding to the phonon momentum).
The small imaginary part in the phonon-assisted dipole moments derives from the exciton–phonon self-energy, see Ref. Lechifflart et al., 2023 . We set its value to 1 meV, which is large enough to remove unphysical divergences and small enough not to affect the final optical spectrum. In this work, we treat the bound exciton as a well-defined quasiparticle and therefore ignore possible overscreening effects.Paleari and Marini (2022); Stefanucci and Perfetto (2025) This approximation has been shown to be excellent in previous studies on BN systems.
II.2 Phonon assisted luminescence
@@
To apply Eqs. (3) and (4) to the light emission process, we use a steady-state approximation to obtain the luminescence via the van Roosbroeck-Shockley (RS) relationVan Roosbroeck and Shockley (1954); Bebb and Williams (1972); Paleari et al. (2019):
In this equation, the oscillator strength of the peak is given by the exciton–phonon coupling matrix elements multiplied by the exciton dipoles . Here and are the energies of the optical and finite-momentum excitons, respectively, while are the phonon energies. The occupation function is
Here, is the temperature-dependent phonon Bose-Einstein occupation function. As can be seen, corresponds to processes of phonon “emission” () — the integer value representing spontaneous phonon emission also at — while corresponds to processes of phonon “absorption” (). Therefore, describes “light” emission by recombining excitons mediated by either “phonon” absorption or emission.
@@
Luminescence is technically an out-of-equilibrium process, but we can assume that for a very low density of excitations and in steady-state conditions, the exciton population can be approximately described by an equilibrium distribution evaluated at an effective temperature . Here, we use the Boltzmann distribution. Experimentally, tends to coincide with the lattice temperature more or less above 100 K, while at very low temperature ( K), may vary between 10–50 K. The parameter is an effective excitonic temperature, used in the Boltzmann distribution to model excitonic occupations. Finally, in a real luminescence calculation, we replace the delta function in Eq. (6) by a Lorentzian, with a smearing set to 0.025 eV.
II.3 Computational details
Ground state properties of the Kohn-Sham bandstructure were obtained using the Quantum-Espresso (QE) codeGiannozzi et al. (2017) using DOJO norm-conserving pseudo-potentialsvan Setten et al. (2018) and PBE-sol exchange correlation functionalPerdew et al. (2008) and a cutoff for the wave-functions of 80 Ry. The phonons were also calculated with the QE code, using a k-grid and a q-grid.
GW and BSE calculations were performed using the Yambo codeSangalli et al. (2019). We used a cutoff of 6 Ha and 80 bands for the dielectric constant, and the same number of bands to expand the Green’s function. In the Bethe-Salpeter equation, we included 3 valence and 3 conduction bands. To speed up convergence with k-point, we used the double-grid method as implemented in Ref. Alliati et al., 2022 with the fine k-point grid of k-points. Exciton–phonon coupling matrix elements were calculated using YamboPyPaleari et al. (2025) and the LetzElPhC libraryLetzElPhC using the same band range of the BSE and the same q-grid.
@@ II.3 Computational details
The symmetry and selection rule analysisNalabothula (2025) was also performed with Yambopy via the spglibTogo et al. (2024) and MolSymGoodlett et al. (2024) packages.
@@
In the calculation of phonon-assisted light absorption, exciton and phonon energies are interpolated on a large q-grid and used in the double-grid version of Eqs. (4) and (6) as described in Appendix A of Ref. Lechifflart et al., 2023. In particular, phonon energies were interpolated using force constants from QEGiannozzi et al. (2017) while excitons were interpolated on the same q-grid using a smooth Fourier interpolation.Pickett et al. (1988) For the calculation of the phonon-assisted dielectric constant we included the lowest 9 excitonic states and we include 20 scattering states in the sum of Eq. 3, that is enough to converge the phonon-assisted absorption/emission, in the luminescence spectrum we consider an excitonic and phononic temperature of 10 K, corresponding to the occupation of the deep minimum of the exciton dispersion.
III Results
In this section, we present the results for the various properties of BN.
III.1 Electronic and phononic bandstructures
@@
In the top panel of Fig. 1, we report the bandstructure of cBN calculated at the KS and G0W0 level. We find the direct band gap at of 8.82 eV and 10.91 eV at DFT and G0W0 levels. The indirect gap between is 4.34 eV and 6.0 eV. These values have to be compared with results from S. Louie et al. Surh et al. (1991) who found 4.3 eV and 6.3 eV for the indirect and 8.6 eV and 11.4 eV for the direct one. G. Satta et al. found 11.4 eV for the direct band gap Satta et al. (2004) and experiments report and 14.5 eV Philipp and Taft (1962). In general we expect our GW correction to be smaller than the one obtained from the S. Louie et al. due to the different plasmon–pole model (for a discussion see Ref. Stankovski et al., 2011), while regarding the comparison with G. Satta et al. two factors can explain the difference: first, we used the experimental lattice constant while they optimized it in DFT, and second, the parameters of our calculations are much more converged. Regarding the experimental result, it must be considered with great uncertainty, given that it is a marginal result with respect to the main aim of that work.
@@
In general, it is well known that in large band gap materials G0W0 tends to underestimate the gap correction and usually self-consistency improves the agreement with the experimentsArtús et al. (2021). However, in this work, we decided to stick with the G0WBSE approximation in order to be consistent with other works in the literature and on the assumption that any possible error in this approximation is of the same order in all BN polymorphs, so that we can compare our results with others on similar materials. We also neglect all energy renormalizations due to electron-phonon interactions, namely quasiparticle band-gap reduction and exciton binding energy increase, which partially compensate. In summary, our results should be interpreted with the understanding that certain physical effects, mostly pertaining to the absolute energy position of the optical spectra, are not taken into account, while they provide a good description of the optical response of cBN as far as coupling strengths and relative energy positions are concerned.
@@
In the bottom panel of Fig. 1, we report the phonon bandstructure calculated with Density Functional Perturbation Theory (DFPT) as implemented in the QE code. With the same approach, we also calculated the deformation potential that enters into the definition of the electron–phonon and then the exciton–phonon matrix elements. We found an LO-TO splitting at of 29 meV and the overall phonon bands are in good agreement with experimental measurements Cai et al. (2007); Eremets et al. (1995) and previous calculationsOhba et al. (2001).
III.2 Optical properties and exciton dispersion
Starting from the quasiparticle bandstructure calculated in the previous section, we solved the Bethe-Salpeter equation (Eq. (2)). Then we calculated the optical absorption of cBN according to Eq. 3. In agreement with previous studies Satta et al. (2004), we found that the optical response of cBN requires a very dense grid to converge. For this reason, we repeated the calculation using the Haydock technique (an iterative solver) combined with a double-grid approachAlliati et al. (2022) to obtain smoother spectra. In Fig. 2 we report the BSE results compared with the independent particle approximation (IPA) starting from the KS bandstructure.
We found that the onset of absorption begins at around 10.5 eV, at much higher energies than those reported in experiments.
In Fig. 2 we have also indicated the first two absorption peaks with letters A and B, the first being a bound state with binding energy of approximately 0.3 eV and the second above the G0W0 gap.
Compared to the IPA, the combination of GW+BSE shifts the spectrum, and several bound excitonic states form.
However, the situation is quite different from that of hBN, where all the spectral weight is concentrated in a few excitonic statesPaleari et al. (2018).
In the cBN case, the spectrum is the sum of many excitonic states with close energies. This is clearly visible when the excitonic dispersion is plotted.
To do this, we solved BSE at finite momentum on a regular q-grid (see Sec. II.3).
In Fig. 3, we report the exciton dispersion along high-symmetry lines. The indirect nature of cBN is clearly visible: in particular, the point (corresponding to the optical limit) does not lie at the bottom of a valley, but instead is one of the highest energy points in the bands.
Here, we have a low-energy subset of 9 exciton states between 10.54 and 10.6 eV. The first absorption peak visible in Fig. 2, denoted as ‘A’, lies in this subset: it is three times degenerate and transforms as the representation of the point group of cBN (only states with this symmetry satisfy the optical selection rules and can be bright).
Furthermore, the LT splitting of this state - verified by solving the BSE at with and without the long-range part of the Coulomb interaction - amounts to 0.05 eV (the longitudinal state jumps from 10.55 to 10.6 eV).
There is also a dark triplet lying 15 meV below the A peak, while the remainder of the 9-state subset is composed by a doubly degenerate state () and a nondegenerate one ().
The minimum of the excitonic dispersion is located at the point at eV, almost 5 eV below the point, consistent with the deep indirect nature of the electronic gap.
Here, the minimum is formed by three states: first, a doubly degenerate state transforming as the representation of the little group of the point, followed 5 meV above by a nondegenerate state.
III.3 Phonon-assisted light absorption/emission
In Fig. 4 we report the phonon-assisted luminescence of cBN calculated using Eq. 6 at 10 K. At this temperature, the lowest-energy exciton () at is occupied, while the state above remains empty. Indeed, the luminescence originates from excitons at the point. We found that the signal is centered around 5.55–5.6 eV, well below that of hBN (from 5.77–5.9 eV). In the same figure, we report the decomposition of the luminescence signal into the contribution of the different phonon modes.
To interpret the luminescence spectrum, we refer to the phonons shown in the lower panel of Fig. 1, highlighted by a vertical red line in the same figure. These are the phonon modes that match the momentum of the lowest excitonic states in Fig. 3. In order to distinguish the different phonons modes we use the same nomenclature as for phonons, as it is clear from the continuous lines of the phonon dispersion which mode we are referring to. Combining all this information we found that phonon-assisted transitions are primarily mediated by transverse phonon branches with a small contribution of longitudinal branches at lower energies.
For the exciton, it is important to emphasize that the selection rules for the phonon-mediated recombination process (which requires a bright exciton as final state, see Eq. 6) allow for coupling with any phonon mode. Instead, in the case of the exciton, the coupling with the mode corresponding to the ‘LA’ branch is forbidden.
Compared to the more famous and better characterized luminescence spectrum of hBNCannuccia et al. (2019); Paleari et al. (2019); Zanfrognini et al. (2023), the spectrum of cBN presents several differences: firstly, it is located at a lower energy; secondly, longitudinal and transverse phonons at are not so separate as in hBN to generate a doublet in luminescence; thirdly, unlike hBN, the phonon modes that contribute most are from the ‘TO’ and ‘TA’ branches, while ‘LA’ and ‘LO’ are low or irrelevant. All these characteristics will make it possible to distinguish it in future experiments on clean samples of cBN.
Finally, in Fig. 5 we report the low energy tail of the dielectric function including the contribution from the exciton–phonon scattering obtained from Eq. 4 interpolated on a double-grid for exciton and phonon energiesLechifflart et al. (2023) at 10 K. We found that coupling with phonons red–shifts the onset from 10.3 to 9.8 eV. Looking at the excitonic dispersion, Fig. 3, one would expect even a lower energy contribution of the epsilon. However, transitions toward low excitons close to have a very weak intensity due to the large denominators that appear in Eq. 5. For this reason, the first visible contribution to the epsilon is only 1 eV below the direct transitions.
IV Conclusions
@@
In this work, we have investigated the optical response of cubic boron nitride (cBN) using first-principles many-body perturbation theory, specifically accounting for the critical role of exciton–phonon coupling.
@@
However, intrinsic exciton–phonon coupling in the pristine system is not enough to reconcile theory with experimental observations. In fact, even if excitons at very low energy are present in cBN, their contribution to the dielectric constant is suppressed by the large denominators in the phonon-assisted dipoles. We expect that the low-energy contribution in the experimental measurement of the dielectric constant of cBN is due to the presence of defects, through direct transition to defect levels or to the activation of transition to finite momentum excitons due to the breaking of translation symmetryTararan et al. (2018).

### v2

# Phonon assisted light absorption and emission in cubic Boron Nitride

@@ Abstract
Cubic boron nitride (cBN) is a wide–bandgap polymorph of boron nitride whose optical response remains only partially understood due to the coexistence of indirect electronic transitions and strong exciton–phonon coupling. Using first-principles Many-Body Perturbation Theory (MBPT), we investigate the optical properties of cBN by combining GW quasiparticle corrections with Bethe-Salpeter equation calculations of excitonic effects. Phonon-assisted absorption and emission processes are explicitly included through the exciton–phonon coupling formalism. We find that phonon–mediated optical transitions provide a dominant contribution to both absorption and luminescence spectra, partially reconciling the discrepancy between the theoretical optical gap ( eV) and experimental emission around 6–7 eV. Our results demonstrate the importance of including exciton–phonon interactions for the correct interpretation of experimental spectra, offering new insights into light emission in wide-bandgap materials.
Boron nitride (BN) exhibits a rich polymorphism, encompassing phases ranging from the common hexagonal (hBN) and cubic (cBN) structures to more complex polytypes.Cazorla and Gould (2019) A significant challenge in studying these materials is the frequent coexistence of different phases and the difficulty in obtaining perfectly pure samples. This complexity underscores the need for precise characterization techniques. While Raman spectroscopy is a widely used characterization tool in materials science, its utility for BN polymorphs is often limited due to the material’s few active Raman modes and its deep UV band gap, which force nonresonant conditions. Consequently, luminescence spectroscopy emerges as a far more powerful and informative probe. The sensitivity of luminescence to electronic structure, defects, and stacking order makes it exceptionally well-suited for characterizing BN nanostructures, where it can reveal details that remain inaccessible to Raman scatteringJaffrennou et al. (2007). For example, luminescence has proven to be a versatile tool, successfully distinguishing between different stacking sequences in the hexagonal phaseZanfrognini et al. (2023) and probing the effects of strain.Lechifflart et al. (2022); Schué (2017)
@@
Here, we present a first-principles investigation of the optical absorption and luminescence of cBN that explicitly includes exciton–phonon coupling within a Many–Body Perturbation Theory (MBPT) framework. We assess the role of indirect excitonic transitions in shaping the optical spectra of cBN and clarify their relevance for the interpretation of experimental observations.
@@
This information is critical, given that cBN is a large-gap material where strongly bound excitons are expected and the coupling between excitons and phonons is known to be crucial in other BN polymorphs.Schué et al. (2019b); Elias et al. (2021b) Furthermore, understanding exciton–phonon coupling in cBN has broader implications, particularly in light of recent studies on small polaron formation in doped cBN systemsShirodkar et al. (2023) and the reported observation of the electron–phonon Fano effect in cBN.Zhu et al. (2023)
The ground-state properties of cubic boron nitride are calculated within density functional theory (DFT), as implemented in the Quantum Espresso (QE) package Giannozzi et al. (2017). The experimental lattice parameter of 3.615 Å reported in Ref. Knittle et al., 1989 is used throughout. Vibrational properties and electron–phonon matrix elements are obtained using density functional perturbation theory (DFPT) Giannozzi et al. (2017). To study the electronic bands of BN, we diagonalized the Kohn-Sham (KS) Hamiltonian and used MBPTMartin et al. (2016) implemented in the Yambo code Sangalli et al. (2019) to correct the KS eigenvalues while keeping the wavefunctions fixed. In this way we obtain the quasiparticle bandstructure within GW approximation, Martin et al. (2016), with the Godby-Needs plasmon–pole model.Stankovski et al. (2011) Starting from the quasiparticle bandstructure we proceeded to calculate the optical response function. To achieve this, we solved the Bethe-Salpeter Equation (BSE) Strinati (1988), which includes both local–field effects and screened electron–hole attraction. The BSE is written on a basis of electron–hole transitions Martin et al. (2016); Gatti and Sottile (2013) and we adopt the Tamm-Dancoff approximation, including therefore only resonant transitions, , and denoting valence and conduction bands, respectively. The matrix elements of the BSE are: where is the quasiparticle energy associated with the above mentioned transition, is the Coulomb potential and is the screened electron–hole interaction.Strinati (1988) After diagonalization of BSE Hamiltonian we obtain the eigenstates and energies of the excitons. The exciton energies and wave-functions at are used to build up the macroscopic dielectric function as:
@@
where are the excitonic dipoles, are the KS dipole matrix elements and is a small broadening term added to mimic the experimental resolution.
@@
For the phonon-assisted absorption/emission, we follow the approach derived in Ref. Lechifflart et al., 2023 with two modifications. First, the phase of the exciton–phonon matrix elements has been fixed using the approach described in Refs. Nalabothula et al., 2025; Nalabothula, 2025. Second, the definition of phonon-assisted transition dipoles has been corrected to be gauge consistent. This corresponds to including also the off-diagonal matrix elements of the exciton-phonon self-energy of Ref. Lechifflart et al., 2023. The first-order correction to the dielectric constant Eq. 1 induced by the coupling with phonons reads:
@@
here, are the phonon-assisted satellite strengths. Eq. 2 describes the satellites generated by the emission or absorption of phonons by an exciton, and they appear as a peak at the energy of the finite-momentum excitons plus or minus one phonon energy . The ‘’ signs refer to phonon emission or absorption processes and , with being the Bose-Einstein occupation function for phonons. The appearance of satellites also induces a renormalization of the static excitonic dielectric constant, a term that we do not consider here because the effect is small, and we investigate the two effects separatelyLechifflart et al. (2023); Marini et al. (2024). The phonon-assisted satellite strengths can be written explicitly as
where are the exciton–phonon coupling matrix elements at momentum Lechifflart et al. (2023); Marini et al. (2024). The small imaginary part in the satellite strengths derives from the exciton–phonon self-energyLechifflart et al. (2023). We set its value to 1 meV, which is large enough to remove unphysical divergences and small enough not to affect the final optical spectrum.
@@
To apply Eqs. (1) and (2) to the light emission process, we use a steady-state approximation to obtain the luminescence via the “excitonic” van Roosbroeck-Shockley (RS) relationVan Roosbroeck and Shockley (1954); Bebb and Williams (1972); Paleari et al. (2019):
@@
In order to model this factor, we suppose a steady-state condition where the exciton population can be approximately described by an equilibrium distribution that we assume to be a Boltzmann function, evaluated at an effective temperature . Finally, in a real luminescence calculation, we replace the delta function in Eq. (4) by a Lorentzian, with a smearing set to 0.025 eV.
Now we briefly present the computational details of the different calculations.
Ground state properties and phonons were obtained using QEGiannozzi et al. (2017) with DOJO pseudopotentialsvan Setten et al. (2018), PBE-sol functionalPerdew et al. (2008), 80 Ry cutoff, and k/q-grids.
GW and BSE calculations were performed using the Yambo codeSangalli et al. (2019). We used a cutoff of 6 Ha and 80 bands for the dielectric constant and the expansion of the Green’s function. In the BSE, we included 3 valence and 3 conduction bands. In agreement with previous studies Satta et al. (2004), we found that the optical response of cBN requires a very dense grid to converge.
In order to accelerate it and obtain smoother spectra, we used the double-grid method Alliati et al. (2022) with the fine k-point grid of k-points in the calculation of the BSE optical absorption spectrum.
Exciton–phonon coupling matrix elements were calculated using YamboPyPaleari et al. (2025) and the LetzElPhC codeLetzElPhC using the same band range and q-grid of the BSE.
@@
The symmetry and selection rule analysisNalabothula et al. (2025) was also performed with Yambopy via the spglibTogo et al. (2024) and MolSymGoodlett et al. (2024) packages.
@@
In the calculation of phonon-assisted light absorption, exciton and phonon energies are interpolated on a large q-grid and used in the double-grid version of Eq. (2) as described in Ref. Lechifflart et al., 2023. In particular, phonon energies were interpolated using force constantsGiannozzi et al. (2017) while excitons were interpolated using a smooth Fourier interpolation.Pickett et al. (1988) For the calculation of the phonon-assisted dielectric constant we included the lowest 9 excitonic states and 100 scattering states in the -sum of Eq. 1, enough to converge the phonon-assisted absorption/emission in the luminescence spectrum.
We consider an excitonic and phononic temperature of 10 K and 300 K, corresponding to the occupation of the deep minimum of the exciton dispersion.
@@
Figure 1 (top panel) shows the bandstructure of cBN calculated at the KS and G0W0 level. We find the direct band gap at of 8.82 / 10.91 eV at DFT / G0W0 levels. The indirect gap between is 4.34 eV / 6.0 eV in the two cases. Surh et al.Surh et al. (1991) found 8.6 / 11.4 eV for the direct and 4.3 / 6.3 eV for the indirect gaps. Satta et al. found 11.4 eV in the G0W0 direct case.Satta et al. (2004). Experiments report and 14.5 eV and for the two gaps Philipp and Taft (1962). In general we expect our GW correction to be smaller than the one obtained from Surh et al. due to the different plasmon–pole modelStankovski et al. (2011), while regarding the comparison with Satta et al., two factors can explain the difference: first, we used the experimental lattice constant while they optimized it in DFT, and second, the parameters of our calculations are much more converged. Regarding the experimental gap result, it must be considered with great uncertainty, given that it is a marginal result with respect to the main aim of that work.
@@
In large band gap materials, G0W0 typically underestimates the gap correction; self-consistency on eigenvalues improves experimental agreementShishkin and Kresse (2007); Artús et al. (2021). However, in this work, we decided to stick with the G0WBSE approximation in order to be consistent with other works in the literature and based on the assumption that any possible error in this approximation is of the same order in all BN polymorphs, so that we can compare our results with others on similar materials. We also neglect all energy renormalizations due to electron-phonon interactions, namely quasiparticle band-gap reduction and exciton binding energy increase, which partially compensate. In summary, our results should be interpreted with the understanding that certain physical effects, mostly pertaining to the absolute energy position of the optical spectra, are not taken into account, while they provide a good description of the optical response of cBN as far as coupling strengths and relative energy positions are concerned.
@@
Figure 1 bottom panel shows the phonon bandstructure calculated with DFPTGiannozzi et al. (2017). With the same approach, we also calculated electron–phonon coupling matrix elements that enter into the definition of the exciton–phonon matrix elements. For reference, we found an LO-TO splitting at of 29 meV and the overall phonon bands are in good agreement with experimental measurements Cai et al. (2007); Eremets et al. (1995) and previous calculationsOhba et al. (2001).
Starting from the quasiparticle bandstructure calculated in the previous section, we solved the BSE. Then we calculated the optical absorption of cBN according to Eq. 1. In Fig. 2 we report the BSE results compared with the independent particle approximation (IPA) starting from the KS bandstructure.
We found that the onset of absorption begins at around 10.5 eV, at much higher energies than those reported in experiments. In Fig. 2 we have also indicated the first two absorption peaks with letters A and B, the first being a bound state with binding energy of approximately 0.3 eV and the second above the G0W0 direct gap. Compared to the IPA, the combination of GW+BSE shifts the spectrum, and several bound excitonic states form (many of them dark). However, the situation is quite different from that of hBN, where all the spectral weight is concentrated in a few excitonic statesPaleari et al. (2018). In the cBN case, the spectrum is the sum of many excitonic states with close energies. This is clearly visible when the excitonic dispersion is plotted. To do this, we solved the BSE at finite momentum on a regular q-grid. In Fig. 3, we report the exciton dispersion along high-symmetry lines. The indirect nature of cBN is clearly visible: in particular, the point does not lie at the bottom of a valley, but instead is one of the highest energy points in the bands. Here, we have a low-energy subset of 9 exciton states between 10.54 and 10.6 eV. The first absorption peak visible in Fig. 2, denoted as ‘A’, lies in this subset: it is three times degenerate and transforms as the representation of the point group of cBN. For reference, the longitudinal–transverse (LT) splitting of this state amounts to 0.05 eV (the longitudinal state jumps from 10.55 to 10.6 eV). There is also a dark triplet lying 15 meV below the A peak, while the remainder of the 9-state subset is composed by a doubly degenerate state () and a nondegenerate one (). The minimum of the excitonic dispersion is located at the point at eV, almost 5 eV below the point, consistent with the deep indirect nature of the electronic gap. Here, the minimum is formed by three states: first, a doubly degenerate state transforming as the representation of the little group of the point, followed 5 meV above by a nondegenerate state.
In Fig. 4 we report the phonon-assisted luminescence of cBN calculated using Eq. 4 at 10 K. At this temperature, the lowest-energy exciton () at is occupied, while the state above remains empty (the luminescence originates from the lowest excitons at the point). We found that the signal is centered around 5.55–5.6 eV, well below that of hBN (from 5.77–5.9 eV). In the same figure, we report the decomposition of the luminescence signal into the contribution of the different phonon modes. To interpret the luminescence spectrum, we refer to the phonons shown in the lower panel of Fig. 1, highlighted by a vertical red line in the same figure. These are the phonon modes that match the momentum of the lowest excitonic states in Fig. 3. In order to distinguish the different phonon modes we use the same nomenclature as for phonons, as it is clear from the continuous lines of the phonon dispersion which mode we are referring to. Combining all this information we found that phonon-assisted transitions are primarily mediated by transverse phonon branches with a small contribution of longitudinal branches at lower energies. For the exciton, it is important to emphasize that the selection rules for the phonon-mediated recombination process allow for coupling with any phonon mode. Instead, in the case of the exciton, the coupling with the mode corresponding to the ‘LA’ branch is forbidden. Compared to the more famous and better characterized luminescence spectrum of hBNCannuccia et al. (2019); Paleari et al. (2019); Zanfrognini et al. (2023), the spectrum of cBN presents several differences: firstly, it is located at a lower energy; secondly, longitudinal and transverse phonons at are not so separate as in hBN to generate a doublet in luminescence; thirdly, unlike hBN, the phonon that contribute most are from the ‘LO’ branches while ‘TO’, ‘LA’ and ‘TA’ are lower in intensity. All these characteristics will make it possible to distinguish it in future experiments on clean samples of cBN. The bottom panel of Fig 4 displays the luminescence at 300 K. At this temperature, the excitons are thermally populated, and their recombinations are primarily mediated by the ‘TA’ and ‘TO’ phonons, each now generating a new luminescence satellite to the higher-energy side of the exciton ones. Interestingly, the ‘TO’ contribution from the A1 exciton is more intense than the corresponding one from the state, pointing to a much larger exciton-phonon coupling for the former that is switched on above K. Finally, in Fig. 5 we report the low energy tail of the dielectric function including the contribution from the exciton–phonon scattering obtained from Eq. 2 interpolated on a double-grid for exciton and phonon energiesLechifflart et al. (2023) at 10 K. We found that coupling with phonons redshifts the onset from 10.3 to 9.6 eV. Looking at the excitonic dispersion, Fig. 3, one would expect even a lower energy contribution of the epsilon. However, transitions toward low excitons close to have a very weak intensity due to the large denominators that appear in Eq. 3. For this reason, the first visible contribution to the epsilon is only 1 eV below the direct transitions.
@@
In conclusion, we have investigated the optical response of cubic boron nitride (cBN) using first-principles many-body perturbation theory, specifically accounting for the critical role of exciton–phonon coupling.
@@
However, intrinsic exciton–phonon coupling in the pristine system is not enough to reconcile theory with experimental observations. In fact, even if excitons at very low energy are present in cBN, their contribution to the dielectric constant is suppressed by the large denominators in the phonon-assisted dipoles. We expect that the low-energy contribution in the experimental measurement of the dielectric constant of cBN is due to the presence of defects, through direct transitions to defect levels or to the activation of transitions to finite momentum excitons due to the breaking of translation symmetryTararan et al. (2018).